# Towards Testing and Evaluating Vision-Language-Action Models for Robotic Manipulation: An Empirical Study

### v1

@@ 4.1. Research Questions
Our empirical study investigates the following research questions to examine the key factors in reliable robotic manipulation: basic performance (RQ1), task complexity (RQ2), perception robustness (RQ3 & RQ 4), and OOD robustness (RQ5).
  • RQ1. How do VLA models perform in popular robotic manipulation tasks? This research question seeks to evaluate the performance of SOTA VLA models in various robotic manipulation tasks. While previous studies (Brohan et al., 2022; Team et al., 2024) typically rely on a few hand-crafted test cases, we use VLATest to generate a large number of test cases. This approach enables a more comprehensive assessment of SOTA VLA models’ performance, providing insights into current challenges and opportunities.
  • RQ2. How does the number of obstacles affect a VLA model’s performance? Intuitively, the presence of more obstacles (i.e., objects unrelated to the assigned manipulation task) increases the complexity of a robotic manipulation task for VLA models. However, it remains unclear whether there is an upper limit to the task complexity that a VLA model can handle effectively. To address this, we conduct controlled experiments to help practitioners understand how VLA models perform as the number of obstacles varies.
  • RQ3. Does the change in lighting conditions affect a VLA model’s performance? When deploying robotics and VLA models in real-world environments, external conditions can vary significantly. Ideally, a VLA model should be robust against different lighting conditions, such as varying illumination intensities. This research aims to offer practitioners practical guidelines on the robustness of VLA models under diverse lighting setups.
  • RQ4. Does the change of camera pose affect a VLA model’s performance? Since VLA models are pre-trained on large-scale vision datasets, they are expected to demonstrate robustness when input images are captured from various angles. This research question seeks to determine whether and to what extent current VLA models maintain robustness when operating with different camera poses.
  • RQ5. How robust do VLA models perform against unseen objects? Objects that were unseen in the robotic demonstration data may present when deploying robotics and VLA models in practical scenarios. It is unclear whether the large-scale pre-training of VLA models can make it generalizable to these unseen objects. Also, there is currently limited understanding of the performance gap between seen-object and unseen-object robotic manipulation tasks. To address this, we investigate this research question by leveraging an external object database to assess the limitations and challenges.
@@ 4.2. Subject VLA Models
To investigate our research questions, we studied four different series of open-source VLA models: RT-1 (Brohan et al., 2022), RT-1-X (Padalkar et al., 2023), Octo (Team et al., 2024), and OpenVLA (Kim et al., 2024). For RT-1 (Brohan et al., 2022), we studied three publicly available variants: RT-1-1k, RT-1-58k, and RT-1-400k  [^1] The number denotes the training steps taken for training an RT-1 model.. We also studied two variants for Octo (Team et al., 2024): Octo-small and Octo-base. Note that, we were unable to study the SOTA VLA model, RT-2 (Brohan et al., 2023), as it has not yet been made publicly available. Thus, we included a total of seven subject VLA models in our empirical study. Table 1 summarizes the characteristics of these models, and we provide more details for each model below.
@@
  • RT-1-X (Padalkar et al., 2023) model was released by Google DeepMind. RT-1-X shares the same model architecture as RT-1. However, RT-1-X was trained on the open-source dataset, Open X-Embodiment (Padalkar et al., 2023), which includes 160k robot demonstrations collected from $2$ different robots.
@@ 4.3. Robotic Manipulation Tasks
We included four different robotic manipulation tasks in our study. Now we briefly introduce each of them. We also refer to the supplementary materials for the video demo of each task.
  • Task 1: Pick up an object. This task requires a VLA model to identify the target object and generate the appropriate control signals to grasp and lift it. To succeed, the robotic must grasp the correct object and lift it at least 0.02 meters for five consecutive frames.
  • Task 2: Move object A to object B. This task requires a VLA model to first identify the source object (A) and then output the corresponding control signals to move it near the target object (B). To succeed, the robotic must move the correct object to within 0.05 meters of the target object.
@@ 4.4. Prompt Templates
For each task, we follow the previous work (Brohan et al., 2022; Padalkar et al., 2023; Kim et al., 2024; Li et al., 2024b) to use the standard prompt template for each task: (1) pick up [object name], (2) move [object name] near [object name], (3) put [object name] on [object name], and (4) put [object name] into [object name].
@@ 4.5. Implementation Details
We conducted all experiments on a server with an AMD 5955WX CPU and two NVIDIA RTX A6000 GPUs. The operating system is 64-bit Ubuntu 20.04 LTS with Python 3.10 and CUDA 12.2. We implemented VLATest with Maniskill2 simulation environments (Gu et al., 2023; Li et al., 2024b). We used two object databases: (1) the default object database in Maniskill2 ($N=18$) in RQ1 ~ RQ4, and (2) YCB object database ($N=56$) in RQ5. The average execution time of one manipulation task was around 19.8 seconds. Our empirical study took about 410 GPU hours in total.
@@ 5.1. RQ1: How Do VLA Models Perform in Popular Robotic Manipulation Tasks?
To investigate this research question, we leveraged VLATest to generate 1,000 scenes for each of the four robotic manipulation tasks by randomly selecting target object(s) and sampling 0 to 3 obstacles. We also randomly assigned poses to these objects. To avoid collision overlaps, we maintained a minimum distance of 0.15 meters between objects during the assignment. The default lighting setups and camera poses were used for this RQ.
Table 2. (RQ1) Performance of seven subject VLA models on different manipulation tasks. The top-1, top-2 and top-3 success rates for each step and task are highlighted, respectively.
@@
All VLA models exhibit subpar performance in the four studied robotic manipulation tasks, particularly in those that require identifying multiple target objects (i.e., Task 2, Task 3, and Task 4). These results suggest that the development of VLA models is still in its early stages, as they are far from being ready for deployment in real-world scenarios.
@@
Current VLA models cannot successfully execute designated task instructions that require multiple steps of action, highlighting the urgent need to improve their capabilities in accurately interpreting natural language prompts.
5.2. RQ2: How Does the Number of Obstacles Affect a VLA Model’s Performance?
To investigate RQ2, we used VLATest to generate 100 scenes for each task with a fixed number of $n$ obstacles, where $n$ was set to 0, 1, 2, 3, and 4, resulting in a total of 500 scenes per task. We used the default lighting conditions and camera poses when investigating this research question.
Figure 4. (RQ2) The number of successful testing scenes at different steps vs. the number of obstacles.
The results are depicted in Fig. 3. In the first two tasks (i.e., Task 1 and Task 2), we observed that the VLA model’s success rate decreased as the number of obstacles increased. The average success rates across different VLA models and scenes dropped from 17.3% to 8.3% and from 8.3% to 1.1% when increasing the number of obstacles from 0 to 4 in Task 1 and Task 2, respectively. However, in Task 3 and Task 4, we did not find a similar pattern. This may be because all VLA models performed poorly even without any obstacles ($n=0$). The average success rates across different VLA models and scenes were 1.2% and 0.7% when $n=0$ and 1.1% and 0.4% when $n=4$ for Task 3 and Task 4, respectively.
The number of obstacles affects the VLA model’s performance, indicating that VLA models become unreliable in more complex environments. When there were four obstacles presented in the scene, the VLA models only passed 8.2%, 2.3%, 1.0%, and 0.4% of the test scenes in Task 1, Task 2, Task 3, and Task 4, respectively.
@@ 5.2. RQ2: How Does the Number of Obstacles Affect a VLA Model’s Performance?
In all four tasks, we found that the success rates for grasping the target object decreased as the number of obstacles increased. These results indicate that when there are multiple obstacles, the VLA models struggle to locate the correct object. Specifically, without any obstacles, the VLA models successfully located the object to manipulate in 199, 196, 137, and 152 out of 700 cases for Task 1, Task 2, Task 3, and Task 4, respectively. However, with four obstacles, the VLA models successfully located the correct object in only 107, 152, 78, and 65 out of 700 cases for the four tasks, respectively. The average success rates of grasping across VLA models dropped to 53.7%, 77.6%, 47.4%, and 42.8% for four tasks, respectively.
With multiple obstacles in the scene, the VLA models’ ability to accurately locate the correct object to manipulate can be significantly affected.
Fig. 5 shows two different testing scenes of Task 1. In both scenes, the VLA model, RT-1-X, was asked to pick up the Pepsi can. When there was no obstacle (Fig. 5(a)), RT-1-X was able to successfully grasp and lift the coke can. However, when there were two other obstacles (i.e., 7up can and Redbull can), RT-1-X failed to locate the Pepsi can and eventually picked up one of the obstacles.
(a) No obstacles
(b) Two obstacles
Figure 5. RT-1-X perform Task 1 with different number of obstacles.
@@ 5.3. RQ3: Does the Change in Lighting Conditions Affect a VLA Model’s Performance?
To answer this research question, we first collected all the successfully executed scenes for each VLA model in RQ1, resulting in a total of 1,434 passed test cases. We then re-executed these test cases three times after randomly increasing or decreasing the lighting intensities ($N=4,302$). Specifically, we randomly sampled a factor $\alpha\in(1,20]$ for increasing or $\alpha\in[1/20,1)$ for decreasing the lighting intensity. This factor was then multiplied by the default lighting intensity used in RQ1. We manually checked the cases when $\alpha$ was set to $20$ and $1/20$ and confirmed that the images captured under such lighting conditions were still recognizable by the humans.
@@
Overall, the VLA models are not sufficiently robust against changes in lighting conditions. Only 61.3% of the test cases could still be successfully executed under mutated lighting conditions. Among the seven VLA models, OpenVLA-7b demonstrated the highest robustness to changes in lighting conditions.
@@
The performance of the VLA models degrades with increasing lighting intensity. When the lighting intensity exceeds 8 times of the default value, the VLA models succeed in only about 40% of the testing scenes that could be passed under default lighting conditions.
Decreasing the lighting intensity ($\alpha<1$). Similar to increasing the lighting intensity, the VLA models’ performance is also affected by even a small perturbation (e.g., $\alpha>0.9$) when decreasing the intensity. The VLA models’ performance degrades with decreasing $\alpha$. However, compared to increasing the lighting intensity, decreasing it has less effects on the models’ performance. When $\alpha<0.2$, the VLA models can still pass around 60% of the testing scenes.
The effect of decreasing the lighting intensity is less significant than increasing it on the VLA models’ performance. Even when the lighting intensity is reduced to 0.15 of its default value, the VLA models can still pass 60% of the test cases.

### v2

@@ 4.1. Research Questions
Our empirical study investigates the following research questions to examine the key factors in reliable robotic manipulation: basic performance (RQ1), task complexity (RQ2), perception robustness (RQ3 & RQ 4), OOD robustness (RQ5), and language model robustness (RQ6).
  • RQ1. How do VLA models perform in popular robotic manipulation tasks?
    This research question seeks to evaluate the performance of SOTA VLA models in various robotic manipulation tasks. While previous studies (Brohan et al., 2022; Team et al., 2024) typically rely on a few hand-crafted test cases, we use VLATest to generate a large number of test cases. This approach enables a more comprehensive assessment of SOTA VLA models’ performance, providing insights into current challenges and opportunities.
  • RQ2. How does the number of confounding objects affect a VLA model’s performance?
    Intuitively, the presence of more confounding objects (i.e., objects unrelated to the assigned manipulation task) increases the complexity of a robotic manipulation task for VLA models. However, it remains unclear whether there is an upper limit to the task complexity that a VLA model can handle effectively. To address this, we conduct controlled experiments to help practitioners understand how VLA models perform as the number of confounding objects varies.
  • RQ3. Does the change in lighting conditions affect a VLA model’s performance?
    When deploying robotics and VLA models in real-world environments, external conditions can vary significantly. Ideally, a VLA model should be robust against different lighting conditions, such as varying illumination intensities. This research aims to offer practitioners practical guidelines on the robustness of VLA models under diverse lighting setups.
  • RQ4. Does the change of camera pose affect a VLA model’s performance?
    Since VLA models are pre-trained on large-scale vision datasets, they are expected to demonstrate robustness when input images are captured from various angles. This research question seeks to determine whether and to what extent current VLA models maintain robustness when operating with different camera poses.
  • RQ5. How robust do VLA models perform against unseen objects?
    Objects that were unseen in the robotic demonstration data may present when deploying robotics and VLA models in practical scenarios. It is unclear whether the large-scale pre-training of VLA models can make it generalizable to these unseen objects. Also, there is currently limited understanding of the performance gap between seen-object and unseen-object robotic manipulation tasks. To address this, we investigate this research question by leveraging an external object database to assess the limitations and challenges.
  • RQ6. How robust do VLA models perform against task instruction mutations?
    The task instruction plays a vital role in the robotic manipulation with VLA models. Task instructions for a VLA model can be paraphrased using different words and sentence structures. Ideally, a VLA model is expected to perform robustly against different mutations of task instructions that carry the same meanings. In this research question, we seek to understand whether VLA models could demonstrate such language model robustness.
@@ 4.2. Subject VLA Models
To investigate our research questions, we studied four different series of open-source VLA models: RT-1 (Brohan et al., 2022), RT-1-X (Padalkar et al., 2023), Octo (Team et al., 2024), and OpenVLA (Kim et al., 2024). For RT-1 (Brohan et al., 2022), we studied three publicly available variants: RT-1-1k, RT-1-58k, and RT-1-400k  [^1] The number denotes the training steps taken for training an RT-1 model.. We also studied two variants for Octo (Team et al., 2024): Octo-small and Octo-base. Note that we were unable to study the SOTA VLA model, RT-2 (Brohan et al., 2023), as it has not yet been made publicly available. Thus, we included a total of seven subject VLA models in our empirical study. Table 1 summarizes the characteristics of these models, and we provide more details for each model below.
@@
Report Performance 92% 73% 30% 71% 71%
(Dataset) RT-1 Paper 6 Skills UR5 WidowX
@@
  • RT-1-X (Padalkar et al., 2023) model was released by Google DeepMind. RT-1-X shares the same model architecture as RT-1. However, RT-1-X was trained on the open-source dataset, Open X-Embodiment (Padalkar et al., 2023), which includes 160k robot demonstrations collected from 22 different robots.
@@ 4.3. Robotic Manipulation Tasks
We included four different robotic manipulation tasks in our study. Now, we briefly introduce each of them. We also refer to the supplementary materials for the video demo of each task.
  • Task 1: Pick up an object. This task requires a VLA model to identify the target object and generate the appropriate control signals to grasp and lift it. To succeed, the robot must grasp the correct object and lift it at least 0.02 meters for five consecutive frames.
  • Task 2: Move object A to object B. This task requires a VLA model to first identify the source object (A) and then output the corresponding control signals to move it near the target object (B). To succeed, the robot must move the correct object to within 0.05 meters of the target object.
@@
For each task, VLATest automatically verifies whether a VLA model has successfully performed it by checking if all corresponding specifications (e.g., the lifted object’s height) are met.
@@ 4.4. Prompt Templates
For each task in RQ1 ~ RQ5, we follow the previous work (Brohan et al., 2022; Padalkar et al., 2023; Kim et al., 2024; Li et al., 2024b) to use the standard prompt template for each task: (1) pick up [object name], (2) move [object name] near [object name], (3) put [object name] on [object name], and (4) put [object name] into [object name]. In RQ6, we use the prompts mutated from these standard templates.
@@ 4.5. Implementation Details
We conducted all experiments on a server with an AMD 5955WX CPU and two NVIDIA RTX A6000 GPUs. The operating system is 64-bit Ubuntu 20.04 LTS with Python 3.10 and CUDA 12.2. We implemented VLATest with Maniskill2 simulation environments (Gu et al., 2023; Li et al., 2024b). We used two object databases: (1) the default object database in Maniskill2 ($N=18$) in RQ1 ~ RQ4 and RQ6, and (2) YCB object database ($N=56$) in RQ5. The average execution time of one manipulation task was around 19.8 seconds. Our empirical study took about 586 GPU hours in total.
@@ 5.1. RQ1: How Do VLA Models Perform in Popular Robotic Manipulation Tasks?
To investigate this research question, we leveraged VLATest to generate 1,000 scenes for each of the four robotic manipulation tasks by randomly selecting target object(s) and sampling 0 to 3 confounding objects. We also randomly assigned poses to these objects. To avoid collision overlaps, we maintained a minimum distance of 0.15 meters between objects during the assignment. The default lighting setups and camera poses were used for this RQ.
Table 2. (RQ1) Performance of seven subject VLA models on different manipulation tasks. The top-1, top-2, and top-3 success rates for each step and task are highlighted, respectively.
@@
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 1, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] All VLA models exhibit subpar performance in the four studied robotic manipulation tasks, particularly in those that require identifying multiple target objects (i.e., Task 2, Task 3, and Task 4). These results suggest that the development of VLA models is still in its early stages, as they are far from being ready for deployment in real-world scenarios.
@@
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 2, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] Current VLA models cannot successfully execute tasks that require multiple steps of action, highlighting the urgent need to improve their capabilities in accurately interpreting task instructions.
Testing Sufficiency. To evaluate the quality of the test scenes generated by VLATest, we further calculated test coverage. However, to the best of our knowledge, no existing coverage metrics are applicable to either VLA models or robotic manipulation. Metrics such as neuron coverage could be computationally intensive to calculate, as VLA models are typically trained on very large datasets with millions or even billions of parameters.
Table 3. Trajectory coverage results with different number of test cases ($n$).
$n$ Task 1 Task 2 Task 3 Task 4
10 0.09 0.09 0.10 0.10
100 0.64 0.66 0.63 0.62
1000 1.00 1.00 1.00 1.00
We followed previous work in autonomous driving to implement trajectory coverage (Hu et al., 2021). We believe this is a pragmatic choice for enabling test coverage measurements. We acknowledge its limitations in Sec. 8. Specifically, we calculated how many novel target object positions were covered w.r.t. the size of the manipulation platform (i.e., the desk). As shown in Table 3, we found that increasing the number of generated test cases $n$ from 10 to 1000 allowed VLATest to achieve 100% coverage across all four tasks, indicating that our generated test scenes are sufficient.
5.2. RQ2: How Does the Number of Confounding Objects Affect a VLA Model’s Performance?
Figure 3. (RQ2) VLA performance vs. the number of confounding objects.
To investigate RQ2, we used VLATest to generate 100 scenes for each task with a fixed number of $n$ confounding objects, where $n$ was set to 0, 1, 2, 3, and 4, resulting in a total of 500 scenes per task. We used the default lighting conditions and camera poses when investigating this RQ.
The results are depicted in Fig. 3. In the first two tasks (i.e., Task 1 and Task 2), we observed that the VLA model’s success rate decreased as the number of confounding objects increased. The average success rates across different VLA models and scenes dropped from 17.3% to 8.3% and from 8.3% to 1.1% when increasing the number of confounding objects from 0 to 4 in Task 1 and Task 2, respectively. However, in Task 3 and Task 4, we did not find a similar pattern. This may be because all VLA models performed poorly even without any confounding objects ($n=0$). The average success rates across different VLA models and scenes were 1.2% and 0.7% when $n=0$ and 1.1% and 0.4% when $n=4$ for Task 3 and Task 4, respectively.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 3, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] The number of confounding objects affects the VLA model’s performance, indicating that VLA models become unreliable in more complex environments. When there were four confounding objects presented in the scene, the VLA models only passed 8.2%, 2.3%, 1.0%, and 0.4% of the test scenes in Task 1, Task 2, Task 3, and Task 4, respectively.
@@ 5.2. RQ2: How Does the Number of Confounding Objects Affect a VLA Model’s Performance?
In all four tasks, we found that the success rates for grasping the target object decreased as the number of confounding objects increased. These results indicate that when there are multiple confounding objects, the VLA models struggle to locate the correct object to manipulate. Specifically, without any confounding objects, the VLA models successfully located the object to manipulate in 199, 196, 137, and 152 out of 700 cases for Task 1, Task 2, Task 3, and Task 4, respectively. However, with four confounding objects, the VLA models successfully located the correct object in only 107, 152, 78, and 65 out of 700 cases for the four tasks, respectively. The average success rates of grasping across VLA models dropped to 53.7%, 77.6%, 47.4%, and 42.8% for four tasks, respectively.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 4, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] With multiple confounding objects in the scene, the VLA models’ ability to accurately locate the correct object to manipulate can be significantly affected.
Figure 4. (RQ2) The number of successful scenes at different steps vs. the number of confounding objects.
(a) No confounding objects
(b) Two confounding objects
Figure 5. RT-1-X perform Task 1 with different number of confounding objects.
Fig. 5 shows two different testing scenes of Task 1. In both scenes, the VLA model, RT-1-X, was asked to pick up the Pepsi can. When there was no confounding object (Fig. 5(a)), RT-1-X successfully grasped and lifted the Pepsi can. However, when there were two confounding objects (i.e., 7 Up can and Red Bull can), RT-1-X failed to locate the Pepsi can and eventually picked up the 7 Up Can.
The Impact of Confounding Objects’ Similarity. In addition to the number of confounding objects, another factor that may impact the performance of VLA models is the similarity between the confounding object(s) and the target object(s). We manually examined our object database and identified two groups of similar objects: (1) pop cans of different brands and (2) cubes of different colors. Then, we categorized our experiment results from RQ2 into two groups: (1) cases where the target object(s) and confounding object(s) were similar and (2) cases where they were dissimilar. We compared the success rates of these two groups under different $n$ across the four tasks. Specifically, we conducted Mann-Whitney U tests, which showed no significant performance difference between the two groups, with $p$-values of $0.443$, $0.614$, $0.657$, and $0.443$ for the four tasks, respectively. Effect sizes are $0.291$, $0.234$, $0.271$, and $0.257$. These results indicate that the similarity between the target objects and the confounding object(s) has little impact on the VLA model’s performance.
@@ 5.3. RQ3: Does the Change in Lighting Conditions Affect a VLA Model’s Performance?
To answer this research question, we first collected all the successfully executed scenes for each VLA model in RQ1, resulting in a total of 1,434 passed test cases. We then re-executed these test cases three times after randomly increasing or decreasing the lighting intensities ($N=4,302$). Specifically, we randomly sampled a factor $\alpha\in(1,20]$ for increasing or $\alpha\in[1/20,1)$ for decreasing the lighting intensity. This factor was then multiplied by the default lighting intensity used in RQ1. We manually checked the cases when $\alpha$ was set to $20$ and $1/20$ and confirmed that the images captured under such lighting conditions were still recognizable by humans.
@@
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 5, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] Overall, the VLA models are not sufficiently robust against lighting changes. Only 61.3% of the test cases could still be successfully executed under mutated lighting conditions. Among the seven VLA models, OpenVLA-7b demonstrated the highest robustness to lighting changes.
@@
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 6, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] The performance of the VLA models degrades with increasing lighting intensity. When the lighting intensity exceeds 8 times the default value, the VLA models succeed in only about 40% of the testing scenes that could be passed under default lighting conditions.
Decreasing the lighting intensity ($\alpha<1$). Similar to increasing the lighting intensity, the VLA models’ performance is also affected by even a small perturbation (e.g., $\alpha>0.9$) when decreasing the intensity. The VLA models’ performance degrades with decreasing $\alpha$. However, compared to increasing the lighting intensity, decreasing it has less significant effects on the models’ performance. When $\alpha<0.2$, the VLA models can still pass around 60% of the testing scenes.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 7, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] The effect of decreasing the lighting intensity is less significant than increasing it on the VLA models’ performance. Even when the lighting intensity is reduced to 0.15 of its default value, the VLA models can still pass 60% of the test cases.